OpenStudy (anonymous):

If apples are bought at 11 for RS 10 & sold at 10 for RS 11, how much will be the gain %?

OpenStudy (anonymous):

cost of apple 10/11 s p of apple = 11/10 11/10 - 10/11 = 21/110 thats the gain per apple use percentage formula to find percentage